### v1

# Uncertainty Matters in Dynamic Gaussian Splatting for Monocular 4D Reconstruction

Abstract
Reconstructing dynamic 3D scenes from monocular input is fundamentally under-constrained, with ambiguities arising from occlusion and extreme novel views. While dynamic Gaussian Splatting offers an efficient representation, vanilla models optimize all Gaussian primitives uniformly, ignoring whether they are well or poorly observed. This limitation leads to motion drifts under occlusion and degraded synthesis when extrapolating to unseen views. We argue that uncertainty matters: Gaussians with recurring observations across views and time act as reliable anchors to guide motion, whereas those with limited visibility are treated as less reliable. To this end, we introduce USplat4D, a novel Uncertainty-aware dynamic Gaussian Splatting framework that propagates reliable motion cues to enhance 4D reconstruction. Our key insight is to estimate time-varying per-Gaussian uncertainty and leverages it to construct a spatio-temporal graph for uncertainty-aware optimization. Experiments on diverse real and synthetic datasets show that explicitly modeling uncertainty consistently improves dynamic Gaussian Splatting models, yielding more stable geometry under occlusion and high-quality synthesis at extreme viewpoints. Project page: https://tamu-visual-ai.github.io/usplat4d/.
Figure 1: We show a challenging case from the DyCheck dataset (Gao et al., 2022), where a person casually rotates a backpack while being captured by a moving monocular camera. The goal is to reconstruct the dynamic object for arbitrary viewpoints and timestamps. The state-of-the-art dynamic Gaussian Splatting methods, e.g., Shape-of-Motion (SoM) (Wang et al., 2025a), struggle on extreme novel views far from the input trajectory, such as opposite-side views (view 1) or large angle offsets (view 2). We propose USplat4D, an Uncertainty-aware dynamic Gaussian Splatting model that produces more accurate and consistent 4D reconstruction. Left: rendered dynamic scene from our model for illustration alongside four sampled RGB inputs. Right: novel view synthesis at two extreme novel viewpoints. Please refer to supplementary video for clearer visual comparison.
1 Introduction
Reconstructing dynamic 3D scenes from monocular input is a fundamental problem across a variety of tasks, including augmented reality, robotics, and human motion analysis (Slavcheva et al., 2017; Li et al., 2023; Newcombe et al., 2015; Joo et al., 2014; Gao et al., 2022). However, despite its wide applicability, monocular dynamic reconstruction remains highly challenging (Liang et al., 2025c), particularly under occlusion and extreme viewpoint changes. Recently, The advent of 3D Gaussian Splatting (Kerbl et al., 2023) has enabled real-time photorealistic rendering and sparked a series of dynamic extensions (Luiten et al., 2024; Lei et al., 2025; Stearns et al., 2024; Liu et al., 2024; Duan et al., 2024; Huang et al., 2024; Yang et al., 2024c; b; Li et al., 2024; Sun et al., 2024b; Duisterhof et al., 2023; Das et al., 2024; Lin et al., 2024; Wu et al., 2025). These methods parameterize motion with shared canonical fields (Wu et al., 2024; Yang et al., 2024c; Liang et al., 2025b; Guo et al., 2024; Lu et al., 2024; Liu et al., 2024; Wan et al., 2024), deformation bases (Wang et al., 2025a; Das et al., 2024; Lin et al., 2024; Li et al., 2024), or direct 4D modeling (Duan et al., 2024; Yang et al., 2024b).
Despite their differences in formulation, existing dynamic Gaussian splatting methods often share a common assumption: motion is optimized uniformly across all Gaussians using 2D supervision such as depth (Yang et al., 2024a), optical flow (Teed & Deng, 2020), and photometric consistency (Doersch et al., 2023). This uniform treatment overlooks that some Gaussians are strongly constrained by recurring observations, while others are only weakly constrained. As a result, motion estimates drift under occlusion and synthesized views degrade at novel viewpoints.
To maintain spatio-temporal consistency, we argue that confidently observed Gaussians should be prioritized and used to guide the optimization of less reliable ones. Consider the example in Figure 1, where a rotating backpack is captured by a moving monocular camera. At any moment, a portion of the surface is self-occluded and invisible. Yet, humans can readily infer their appearance and motion by recalling previously observed surfaces and extrapolating with temporal continuity. Such ability anchors on the most reliable parts of the backpack, i.e., those clearly observed from other viewpoints and timestamps. This suggests a key principle: when observations are partial, reconstruction should be guided by confident cues and propagated structurally to uncertain regions.
Building on this insight, we propose USplat4D, a novel Uncertainty-aware dynamic Gaussian Splatting framework for monocular 4D reconstruction. We first introduce a principled method to estimate time-varying uncertainty for each Gaussian, capturing how reliably it is constrained by recurring observations. This uncertainty then guides the selection of anchor Gaussians and propagate motion across space and time. To realize this principle, we organize Gaussians into a spatio-temporal graph, where uncertainty determines node importance, edge construction, and adaptive loss weighting. The goal of the uncertainty-aware graph optimization is to ensure that confident parts of the scene dynamically guide the reconstruction of the rest, even in occluded or unseen views.
We validate our approach on various real and synthetic datasets on monocular 4D reconstruction. We show that explicitly leveraging uncertainty significantly enhances both motion tracking and novel view synthesis, with particularly strong gains under extreme viewpoints. Our framework, including uncertainty estimation, graph construction, and adaptive training, is model-agnostic and can be integrated into existing dynamic Gaussian splatting pipelines that parameterize per-Gaussian motion. Overall, USplat4D introduces a principled way to model uncertainty in dynamic Gaussian Splatting, yielding more stable motion estimates under occlusion and high-quality extreme view synthesis.
2 Related Work
Dynamic Gaussian splatting. Recent advances in dynamic Gaussian Splatting have enabled monocular 4D reconstruction via per-Gaussian deformation or canonical motion modeling (Liang et al., 2025c; Yang et al., 2024c; b; Wu et al., 2024; Li et al., 2024) or high-fidelity dynamic scene reconstruction from multi-view inputs (Luiten et al., 2024; Wang et al., 2025b). To reconstruct 4D GS from the monocular video, methods such as SoM (Wang et al., 2025a), MoSca (Lei et al., 2025), Marbles (Stearns et al., 2024), and 4D-Rotor (Duan et al., 2024) use low-rank motion bases to regularize deformation, while others model canonical flows (Liang et al., 2025b; Liu et al., 2024). Although they demonstrate high-fidelity rendering on near-input validation views, they do not explicitly model the motion behind occluders or identify reliable Gaussians for motion guidance. MoSca (Lei et al., 2025) introduces a soft motion score but lacks structured propagation. In contrast, USplat4D selects high-confidence Gaussians and constructs an uncertainty-aware motion graph to propagate motion through spatio-temporally coherent connections, improving robustness in occluded regions and enabling localized refinement beyond low-rank modeling (Kim et al., 2024; Huang et al., 2024).
Uncertainty estimation in scene reconstruction. Uncertainty modeling has been widely explored in neural rendering (Li et al., 2022; Pan et al., 2022; Shen et al., 2021; 2022; Kim et al., 2022; Zhan et al., 2022; Yan et al., 2023; Lee et al., 2022; Sünderhauf et al., 2023), particularly to improve robustness under occlusion, sparse views, and ambiguity. To avoid overfitting on reconstructing the static scene, SE-GS (Zhao et al., 2025) designs an uncertainty-aware perturbing strategy by estimating the self-ensembling uncertainty. In dynamic settings, uncertainty has been used to smooth motion or reweight gradients (Kim et al., 2024), but typically as an auxiliary signal decoupled from the underlying motion representation. In contrast, USplat4D treats uncertainty as a central modeling component. We estimate confidence of each Gaussian and use it to guide key node selection, edge construction, and loss weighting in a spatio-temporal graph. This allows high-confidence Gaussians to guide motion propagation while reducing the influence of uncertain regions. To the best of our knowledge, this is among the first attempts and analyses to model the uncertainty and directly integrate it into graph-based motion modeling for dynamic reconstruction.
3 Preliminary
We first review dynamic Gaussian Splatting and its learning objective to establish the notation for Section 4. Our method is model-agnostic, building on this standard formulation and applicable to a wide range of dynamic Gaussian Splatting variants (Lei et al., 2025; Wang et al., 2025a).
Dynamic 3D Gaussians. Vanilla dynamic Gaussian Splatting (Luiten et al., 2024) represents a scene with a set of time-varying 3D Gaussians. Formally, the state of a Gaussian at time $t$ is defined as
$\mathbf{G}_{t}=(\mathbf{p}_{t},\mathbf{q}_{t},\mathbf{s},\alpha,\mathbf{c}),$ (1)
where $\mathbf{p}_{t}\in\mathbb{R}^{3}$ denotes the position at time $t$, $\mathbf{q}_{t}\in\mathbb{R}^{4}$ the quaternion rotation, $\mathbf{s}\in\mathbb{R}^{3}$ the scale, $\alpha\in\mathbb{R}$ the opacity, and $\mathbf{c}\in\mathbb{R}^{N_{c}}$ the color coefficients (e.g., spherical harmonics or RGB), $N_{c}$ the color dimension. The trajectory of a Gaussian is then given by the sequence $\{\mathbf{G}_{t}\}_{t=1}^{T}$, where $T$ is the number of frames. While early extensions introduce time-varying color (Yang et al., 2024b), this hinders 3D motion tracking. Recent methods, such as SoM (Wang et al., 2025a) and MoSca (Lei et al., 2025), restrict the color space to stabilize motion estimation and enable reliable tracking.
Learning objectives. To optimize the 4D Gaussian field, existing methods minimize a combination of losses. A photometric reconstruction loss enforces consistency between rendered and ground-truth images, while motion-locality losses regularize the temporal evolution of Gaussians. These locality terms include isometry, rigidity, relative rotation, velocity, and acceleration constraints (Lei et al., 2025; Huang et al., 2024), which shrink the large motion search space and stabilize optimization.
Limitations. Although effective near input views, these objectives remain fragile under occlusion and extreme novel viewpoints, as they rely heavily on unstable 2D priors such as depth, optical flow, or photometric consistency. As a result, reconstructions often drift over time and lose geometric consistency across different views. To overcome this challenge, we introduce a dynamic uncertainty model that explicitly encodes the reliability of each Gaussian over time and forms the foundation of our framework. In particular, Gaussians are partitioned into key and non-key nodes and connected through an uncertainty-weighted graph, which enforces spatio–temporal consistency.
Figure 2: Overview of the proposed USplat4D. We first estimate time-varying uncertainty for each Gaussian (Section 4.1). We then leverage these uncertainties to select reliable Gaussians as key nodes, while others are treated as non-key nodes for graph construction (Section 4.2). Finally, we optimize the spatio-temporal graph with uncertainty-weighted losses, yielding consistent 4D Gaussians (Section 4.3). The right column shows that our approach significantly improves novel view renderings compared to vanilla optimization.
4 Uncertainty-Aware Dynamic Gaussian Splatting
Overview. Given a monocular video, we begin by building on vanilla dynamic Gaussian Splatting models to estimate a time-varying uncertainty score for each Gaussian, explicitly capturing its reliability across frames (Section 4.1). These uncertainty scores then guide the construction of an uncertainty-weighted graph that systematically organizes Gaussians into key and non-key nodes (Section 4.2). The resulting graph subsequently drives an optimization process that propagates reliable motion cues to uncertain regions, thereby refining both motion estimation and rendering quality of the dynamic scene (Section 4.3). An overview of the entire pipeline is illustrated in Figure 2.
4.1 Dynamic Uncertainty Estimation
Existing dynamic Gaussian Splatting optimizes all primitives uniformly, even though some are well observed across time while others remain ambiguous. This causes drift under occlusion and instability at extreme viewpoints. We therefore assign each Gaussian a time-varying uncertainty to estimate its reliability per frame and use it to guide optimization in a model-agnostic way.
Per-Gaussian scalar uncertainty. A straightforward way to capture reliability is to assign each Gaussian $i$ a scalar uncertainty $u_{i,t}$ at every frame $t$. Intuitively, if a primitive is frequently and clearly observed, its uncertainty should be small; if it is rarely visible or weakly constrained, its uncertainty should be large. The photometric loss is defined as
$\mathcal{L}_{2,t}=\sum\nolimits_{h\in\Omega}\|\bar{C}_{t}^{h}-C_{t}^{h}\|_{2}^{2},\quad\text{where}\quad C_{t}^{h}$ $=\sum\nolimits_{i=1}^{N_{g}}T_{i,t}^{h}\alpha_{i}\,c_{i}.$ (2)
Here, $\Omega$ is the pixel index set, $\bar{C}_{t}^{h}$ and $C_{t}^{h}$ denote the ground-truth and rendered colors at pixel $h$, and ${c}_{i}$ is the color parameter of Gaussian $i$ at frame $t$. The rendered pixel color is obtained by $\alpha$-blending, where the blending weight is given by $T_{i,t}^{h}\alpha_{i}$, with $T_{i,t}^{h}$ the transmittance of Gaussian $i$ at pixel $h$, $\alpha_{i}$ its opacity, and $N_{g}$ is the number of Gaussians. By differentiating $\mathcal{L}_{2,t}$ with respect to ${c}_{i}$ and applying the local minimum assumption, we obtain the closed-form variance estimate (please see Appendix A.1 for detailed derivation):
$\sigma_{i,t}^{2}=\left(\sum\nolimits_{h\in\Omega_{i,t}}(T_{i,t}^{h}\alpha_{i})^{2}\right)^{-1},$ (3)
where $\Omega_{i,t}\subseteq\Omega_{t}$ denotes the set of pixels contributed by that Gaussian. We thus take this variance as the scalar uncertainty, i.e., $u_{i,t}:=\sigma_{i,t}^{2}$. However, the local minimum assumption may not hold everywhere. To account for unconverged pixels, we introduce an indicator function to test per-pixel convergence:
$\mathcal{I}(h)=\left\{\begin{array}[]{ll}1&\quad\textrm{if }\|\bar{C}_{t}^{h}-C_{t}^{h}\|_{1}<\eta_{c},\\ 0&\quad\textrm{otherwise},\end{array}\right$. (6)
where $\eta_{c}>0$ is a color-error threshold. For Gaussian $i$ at frame $t$, we define the aggregate indicator $\mathcal{I}_{i,t}=\prod_{h\in\Omega_{i,t}}\mathcal{I}_{t}(h),$ which equals $1$ only if all covered pixels are convergent. If $\mathcal{I}_{i,t}=0$, we assign a large constant $\phi$ to indicate high uncertainty. Therefore, the final scalar uncertainty is:
$u_{i,t}=\mathcal{I}_{i,t}\,\sigma_{i,t}^{2}+(1-\mathcal{I}_{i,t})\,\phi.$ (7)
This design jointly constraint strength with convergence status: Gaussians that are well supported by observations receive low $u_{i,t}$, reflecting high reliability, while unreliable ones are assigned high values, which enables trustworthy primitives to guide ambiguous ones during optimization.
From scalar to depth-aware uncertainty. While the scalar definition is intuitive, it implicitly assumes that uncertainty is isotropic in 3D space. This is problematic in monocular settings, where depth is much less reliable than image-plane coordinates. A uniform scalar tends to be over-confident along the camera axis, leading to geometric distortion. To address this, we propagate image-space errors into 3D and represent each Gaussian by an anisotropic uncertainty matrix:
$\mathbf{U}_{i,t}=\mathbf{R}_{wc}\,\mathbf{U}_{c}\,\mathbf{R}_{wc}^{\mathsf{T}},\quad\text{where}\quad\mathbf{U}_{c}=\text{diag}(r_{x}u_{i,t},\,r_{y}u_{i,t},\,r_{z}u_{i,t}).$ (8)
Here, $\mathbf{R}_{wc}$ is the camera-to-world rotation and $r_{x},r_{y},r_{z}$ are axis-aligned scaling factors. Note that only rotation is required to propagate uncertainty, since translation does not affect covariance. This transforms 2D uncertainty into axis-aligned 3D uncertainty, incorporating both the camera pose and the directional sensitivity of depth. A typical example is the “Camel” sequence in Figure 2 (also see the supplementary video): without depth-aware uncertainty, the camel’s body shrinks unnaturally, whereas our formulation preserves its correct shape.
4.2 Uncertainty-encoded Graph Construction
Per-Gaussian uncertainty in Equation 8 provides a local measure of reliability, but treating Gaussians independently cannot guarantee spatio–temporal consistency. Neighboring primitives often share correlated motion, and reliable ones should anchor the optimization of uncertain ones. Prior graph-based methods (Huang et al., 2024; Lei et al., 2025) attempt to capture this correlation, e.g., MoSca (Lei et al., 2025) introduces a 3D lifting graph. In contrast, we build the graph directly on uncertainty: Gaussians are ranked by reliability and partitioned into key and non-key nodes, so that stable primitives drive motion propagation while ambiguous ones are regularized. To realize this, we design an uncertainty-aware graph that encodes reliability in both node selection and edge connectivity, providing the foundation for the optimization described in the following sections.
Graph definition. We represent the scene with a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where each node $i\in\mathcal{V}$ corresponds to a Gaussian $\mathbf{G}_{i}$ and edges $(i,j)\in\mathcal{E}$ encode spatial affinity and motion similarity. Crucially, nodes are partitioned into a small key set $\mathcal{V}_{k}$ and a large non-key set $\mathcal{V}_{n}$ according to their uncertainties $\{u_{i,t}\}$ from Section 4.1. We define key nodes as stable Gaussians that carry strong motion cues across time and views, while non-key nodes inherit motion from their key neighbors.
Key node selection. Our key node selection operates in 3D space. We employ the following two-stage strategy that balances spatial coverage and temporal stability:
  1. 1.
[…]
  2. 2.
    Thresholding by significant period. For each Gaussian candidate, we compute its significant period, defined as the number of frames where its uncertainty stays below a threshold. We retain only those with a significant period of at least 5 frames, ensuring that key nodes have sufficient temporal support to contribute reliably to motion estimation. Candidates with insufficient temporal coverage tend to destabilize graph optimization and yield under-constrained solutions.
Edge construction. We construct edges separately for key and non-key nodes, as their roles differ: key nodes provide structural anchors for motion propagation, while non-key nodes interpolate appearance locally. The key graph captures long-range geometric and motion dependencies (e.g., both ends of a limb moving coherently), which existing distance-based heuristics such as local kkNN (Huang et al., 2024) or global min–max distances (Lei et al., 2025) cannot robustly handle.
To address this, we adopt an Uncertainty-Aware kkNN (UA-kkNN). For a key node $i$, we select neighbors only among other low-uncertainty key nodes, evaluated at its most reliable frame $\hat{t}=\arg\min_{t}\{u_{i,t}\}$, and measure distances with uncertainty weighting to favor trustworthy connections:
$\mathcal{E}_{i}=\text{$k$NN}_{j\in\mathcal{V}_{k}\setminus\{i\}}\Big(\|\mathbf{p}_{i,\hat{t}}-\mathbf{p}_{j,\hat{t}}\|_{(\mathbf{U}_{w,\hat{t},i}+\mathbf{U}_{w,\hat{t},j})}\Big).$ (9)
Here, the Mahalanobis metric up-weights directions of high uncertainty, so edges are formed between nodes that are both spatially close and reliable. As will be shown in Section 4.3, these edges are further pruned by the key graph loss for additional robustness and to prevent spurious long-range connections. For a non-key node $i$, we assign it to its closest key node across the sequence:
$j={\arg\min}_{l\in\mathcal{V}_{k}}\sum\nolimits_{t=0}^{T-1}\|\mathbf{p}_{i,t}-\mathbf{p}_{l,t}\|_{(\mathbf{U}_{w,t,i}+\mathbf{U}_{w,t,l})}$ (10)
and connect $\mathcal{E}_{i}=\mathcal{E}_{j}\cup\{j\}$ and $j$ is the closet key nodes. Intuitively, each uncertain non-key node is attached to the most reliable key node that stays close to it over time, so its motion can be regularized by stable anchors for consistency. In both cases, uncertainty-aware kkNN ensures edges are anchored by reliable nodes, promoting stable motion propagation and preventing drift from uncertain regions.
4.3 Uncertainty-Aware Optimization
Vanilla optimization (Section 3) of dynamic Gaussians often fail under occlusion or extreme viewpoints. This leads unreliable primitives to drift, since they are optimized as strongly as reliable ones. To address this, we incorporate uncertainty into the objective: key nodes with stable observations serve as anchors, while non-key nodes are regularized more softly through interpolation. We design separate objectives for key and non-key nodes, then unify them in a total loss.
Key node loss. Key nodes are low-uncertainty Gaussians that anchor motion. We encourage them to stay close to their pre-optimized positions:
$\mathcal{L}^{\textrm{key}}=\sum\nolimits_{t=0}^{T-1}\sum\nolimits_{i\in\mathcal{V}_{k}}\|\mathbf{p}_{i,t}-\mathbf{p}_{i,t}^{\mathrm{o}}\|_{\mathbf{U}_{w,t,i}^{-1}}+\mathcal{L}^{\textrm{motion,key}},$ (11)
where $\mathbf{U}_{w,t,i}$ down-weights directions of high uncertainty, ensuring motion is corrected mainly along reliable axes. Superscript $\mathrm{o}$ denotes the parameters from the pretrained Gaussian Splatting model, which serves as initialization before our uncertainty-aware optimization. The key node motion loss $\mathcal{L}^{\textrm{motion,key}}$ regularizes the temporal evolution of Gaussians by isometry, rigidity, rotation, velocity, and acceleration constraints, which is discussed in the Appendix A.2 in detail.
Non-key node loss. Non-key nodes are interpolated from nearby key nodes using Dual Quaternion Blending (DQB) (Kavan et al., 2007), which provides smooth motion by blending their neighbors:
$\left(\mathbf{p}_{i,t}^{\textrm{DQB}},\mathbf{q}_{i,t}^{\textrm{DQB}}\right)=\text{DQB}\Big(\{(w_{ij},\mathbf{T}_{j,t})\}_{j\in\mathcal{E}_{i}}\Big),$ (12)
where $w_{ij}$ are normalized edge weights for blending, $\mathbf{T}_{j,t}\in\mathbb{SE}(3)$ is the transform of key node $j$, and $\mathbf{p}_{i,t}^{\textrm{DQB}}$ and $\mathbf{q}_{i,t}^{\textrm{DQB}}$ are the position and rotation of Gaussian $i$ at time $t$ acquired by DQB, respectively. We then regularize non-key nodes to both their initialization and interpolated trajectory:
$\mathcal{L}^{\textrm{non-key}}=\sum\nolimits_{t=0}^{T-1}\sum\nolimits_{i\in\mathcal{V}}\|\mathbf{p}_{i,t}-\mathbf{p}_{i,t}^{\mathrm{o}}\|_{\mathbf{U}_{w,t,i}^{-1}}+\sum\nolimits_{t=0}^{T-1}\sum\nolimits_{i\in\mathcal{V}}\|\mathbf{p}_{i,t}-\mathbf{p}_{i,t}^{\textrm{DQB}}\|_{\mathbf{U}_{w,t,i}^{-1}}+\mathcal{L}^{\textrm{motion,non-key}}.$ (13)
Here, $\mathcal{L}^{\textrm{motion,non-key}}$ is the non-key node motion loss, which will also be discussed in Appendix A.2. This non-key loss keeps non-key nodes close to their pretrained states while aligning them with motions propagated from reliable key nodes, preventing drift while ensuring coherence.
Total loss. The final objective combines the key and non-key node losses with the photometric loss:
$\mathcal{L}^{\textrm{total}}=\mathcal{L}^{\textrm{rgb}}+\mathcal{L}^{\textrm{key}}+\mathcal{L}^{\textrm{non-key}}.$ (14)
Overall, the uncertainty in our framework serves three purposes: (1) re-weighting deviations of key nodes, (2) guiding the interpolation of non-key nodes, and (3) balancing their influence in the total loss. Our design mitigates drift under occlusion and maintains geometric consistency at novel views.
5 Experiments
Figure 3: Qualitative results on validation views of the DyCheck dataset (Gao et al., 2022). We show comparisons with two strong baselines, SoM (Wang et al., 2025a) and MoSca (Lei et al., 2025). USplat4D improves visual quality and better preserves geometry (e.g., arms in “Spin”, hands in “Space-out”) and pose (e.g., “Windmill”). Please zoom in for details. See supplementary for more examples and other baselines.
Figure 4: Qualitative results on extreme novel views from DyCheck. Unlike Figure 3, these manually sampled extreme views (red cameras) lack ground-truth. USplat4D better preserves fine structures (e.g., the dog’s head in “haru-sit”) and occluded regions (e.g., the hand in “spin”) under large viewpoint shifts.
Table 1: Quantitative results on DyCheck. We report results on 5 scenes at $1\times$ resolution and 7 scenes at $2\times$ resolution, following existing protocols. USplat4D consistently outperforms state-of-the-art Gaussian Splatting based methods. See Figure 3 for qualitative results on validation views and Figure 4 for extreme views.
[…]
5 scenes
1 $\times$ resolution
SC-GS (Huang et al., 2024) 14.13 0.477 0.49
Deformable 3DGS (Yang et al., 2024c) 11.92 0.490 0.66
4DGS (Wu et al., 2024) 13.42 0.490 0.56
MoDec-GS (Kwak et al., 2025) 15.01 0.493 0.44
MoBlender (Zhang et al., 2025) 16.79 0.650 0.37
SoM (Wang et al., 2025a) 16.72 0.630 0.45
USplat4D (ours) 16.85 0.650 0.38
7 scenes
2 $\times$ resolution
Dynamic Gaussians (Luiten et al., 2024)   7.29 0.69
4DGS (Wu et al., 2024) 13.64 0.43
Gaussian Marbles (Stearns et al., 2024) 16.72 0.41
MoSca (Lei et al., 2025) 19.32 0.706 0.26
USplat4D (ours) 19.63 0.716 0.25
5.1 Datasets and Baselines
Datasets. (1) DyCheck (Gao et al., 2022): We follow prior works (Wang et al., 2025a; Lei et al., 2025; Stearns et al., 2024) to evaluate on 7 scenes with validation views. Since these validation views are near the input views, we additionally sample extreme novel views for qualitative analysis. (2) DAVIS (Perazzi et al., 2016): To test generalization across different scenarios, we qualitatively evaluate on challenging monocular videos using DAVIS dataset, which cover non-rigid motion, occlusions, and complex dynamics. (3) Objaverse (Deitke et al., 2023): We construct a synthetic benchmark [^3] Existing benchmarks lack ground-truth for extreme novel views, making systematic evaluation difficult. We therefore build a synthetic benchmark, where strong results support our generalization claims (see experiments). Note that such synthetic setups are widely used to stress-test models (Yao et al., 2025; Liang et al., 2024). from Objaverse by selecting 6 challenging articulated objects with diverse textures and motions. Specifically, we set cameras to follow a circular trajectory of 121 frames (3 per step) around each object, always facing toward it. For validation, we render views at horizontal angular offsets of 30° increments from 30° to 330°, each with a fixed elevation of 35°, in order to capture the object’s 3D structure from diverse viewpoints. Please see Appendix C.4 for additional details.
Baselines. We compare USplat4D with state-of-the-art dynamic Gaussian Splatting methods. We use SoM (Wang et al., 2025a) and MoSca (Lei et al., 2025) as base models. SoM is widely adopted, and MoSca represents the current state of the art. Please see Appendix B for additional details.
Figure 5: Qualitative results on extreme novel views from DAVIS. For each case, we show an input-view rendering and compare the baseline (SoM (Wang et al., 2025a) or MoSca (Lei et al., 2025)) with our USplat4D on an extreme novel view (red). USplat4D yields clearer reconstructions under challenging conditions.
Table 2: Results on the Objaverse dataset. We evaluate novel view synthesis across increasing horizontal angular ranges: $(0^{\circ},60^{\circ}]$, $(60^{\circ},120^{\circ}]$, and $(120^{\circ},180^{\circ}]$. USplat4D consistently improves over SoM (Wang et al., 2025a) and MoSca (Lei et al., 2025), with gains most pronounced at larger viewpoint shifts.
Method View Range $(0^{\circ},60^{\circ}]$ View Range $(60^{\circ},120^{\circ}]$ View Range $(120^{\circ},180^{\circ}]$
[…]
SoM 16.09 0.860 0.31 15.58 0.854 0.32 16.45 0.858 0.31
USplat4D 16.63.55 0.866.007 0.27.03 16.57.09 0.868.014 0.27.05 17.03.58 0.872.014 0.26.05
MoSca 16.18 0.881 0.24 15.74 0.875 0.25 15.89 0.876 0.25
USplat4D 16.22.04 0.885.004 0.22.02 15.98.24 0.884.009 0.23.02 16.31.42 0.886.011 0.21.03
Figure 6: Qualitative results on Objaverse. Each case shows a 3D rendering from an input view and a comparison between the baseline (SoM (Wang et al., 2025a) or MoSca (Lei et al., 2025)) and our USplat4D at an extreme novel view (red). Please see the supplementary video for clearer visualization.
5.2 Main Results
Results on DyCheck. Table 1 reports results on DyCheck validation views, which are close to the input trajectories. USplat4D consistently outperforms baselines across nearly all metrics; qualitative examples are shown in Figure 3. However, these validation views remain relatively easy. The more significant improvements emerge under extreme novel viewpoints (Figure 4), which are not included in Table 1 but are essential for assessing robustness under severe viewpoint shifts. Please see Appendix C.1 for extra baselines, and Appendix C.2 and C.3 for tracking and per-scene results.
Qualitative results of extreme novel view synthesis on DAVIS. We evaluate on selected monocular videos from the DAVIS dataset, which include fast motion, deformation, and self-occlusion. As ground-truth geometry is unavailable, we focus on qualitative comparisons. As shown in Figure 5, USplat4D yields more plausible geometry and coherent reconstructions under extreme viewpoint shifts, where the baseline often exhibits distortion, blur, or artifacts.
Results on Objaverse. Table 2 shows that USplat4D consistently surpasses SoM (Wang et al., 2025a) and MoSca (Lei et al., 2025), with gains most pronounced under large viewpoint shifts ($120^{\circ}$$180^{\circ}$). Figure 6 confirms these improvements qualitatively: our method preserves geometry and textures under extreme novel views, where baselines often blur or collapse. We include more details in Appendix C.5, including dataset setups and per-scene results.
5.3 Ablation and Analysis
We conduct ablation studies to assess key design choices in USplat4D, using MoSca (Lei et al., 2025) as the base model and evaluating on the DyCheck validation set (Gao et al., 2022).
Table 3: Ablation study on uncertainty usage. We assess the impact of uncertainty estimation by removing it from key components in USplat4D individually.
[…]
USplat4D (full model) 19.63 0.716 0.25
(a) w/o key node uncertainty 18.86 0.688 0.28
(b) w/o UA-kkNN 19.50 0.711 0.26
(c) w/o loss weighting 19.08 0.681 0.25
Table 3 shows that uncertainty is essential across key components of our framework. (a) When removed from key node selection and replaced with uniform 2D sampling, the graph fails to emphasize well-constrained Gaussians, leading to unstable anchors and degraded propagation. (b) Replacing UA-kkNN with distance-only kkNN weakens graph connectivity, as edges disregard node reliability and mistakenly connect unstable primitives. (c) Excluding uncertainty weighting from the training loss (i.e., applying the DQB loss without $\mathbf{U}$) reduces PSNR/SSIM, since unreliable Gaussians are updated as aggressively as reliable ones, causing drift across frames. Please see Appendix D additional discussion and other analysis.
6 Discussion and Conclusion
We introduce a novel dynamic Gaussian Splatting framework USplat4D, showing that uncertainty matters in dynamic Gaussian splatting. We propose time-varying uncertainty estimation method and built an uncertainty-guided graph, systematically demonstrating improved geometric consistency and rendering quality under extreme views. While our method still inherits errors from the underlying Gaussian splatting model, we hope that our findings could encourage further work on leveraging uncertainty to advance robust 4D reconstruction. Please see appendix for more discussion.
Appendix
This appendix provides additional theoretical derivations, implementation details, experimental results, ablation studies, and discussions on broader impacts. A quick outline is as follows:
  • Section A: Additional details on our proposed USplat4D.
    • Section A.1: Derivation of the uncertainty model.
    • Section A.2: Details of the training losses.
  • Section B: Implementation details, including adaptation of base models and training schedules.
    • Section B.1: Adaptation of SoM and MoSca.
    • Section B.2: Training USplat4D Model.
  • Section C: Additional experimental results.
    • Section C.1: Comparison with additional NeRF-based methods on the DyCheck dataset.
    • Section C.2: Tracking results on the DyCheck dataset.
    • Section C.3: Per-scene results on the DyCheck dataset.
    • Section C.4: Data curation for the Objaverse dataset.
    • Section C.5: Per-scene results on the Objaverse dataset.
    • Section C.6: Results on the NVIDIA dataset.
  • Section D: Additional ablation studies and analyses.
    • Section D.1: Ablation study on key / non-key ratio.
    • Section D.2: Analysis on key node selection.
    • Section D.3: Analysis on runtime.
    • Section D.4: Failure cases.
  • Section E: Discussion of the social and broader impact of our work.
Appendix A Additional Details on USplat4D
A.1 Derivation of Uncertainty (Section 4.1 of the Main Paper)
To derive our uncertainty model, we begin by analyzing the color blending function used in volumetric rendering. To simplify the notation, we omit time $t$. The rendered pixel color $C$ along a ray is computed as a weighted sum of the color contributions from all Gaussians:
$C^{h}=\sum_{i=1}^{N_{g}}T_{i}^{h}\alpha_{i}c_{i}:=\sum_{i=1}^{N_{g}}v_{i}^{h}c_{i},$ (S1)
where $i$ is the index of Gaussians, $T_{i}$ is the accumulated transmittance up to the $i$-th Gaussian, $\alpha_{i}$ is the opacity, $c_{i}$ is the color, and $v_{i}:=T_{i}\alpha_{i}$ denotes the blending weight. To estimate the color uncertainty, we derive a closed-form expression for each Gaussian’s color via maximum likelihood estimation (MLE) under an RGB $\ell_{2}$ loss:
$\mathcal{L}_{2}=\sum_{h\in\Omega}\|\bar{C}^{h}-C^{h}\|_{2}^{2}$ (S2)
Here, $h\in\Omega$ indexes the set of all pixels, $\bar{C}^{h}$ is the ground-truth color, and $C^{h}$ is the color rendered using the blending model in Eq. S1.
To find the optimal color $c_{k}$ for the $k$-th Gaussian, we compute the gradient of the loss with respect to $c_{k}$ and obtain:
$\left.\frac{\partial\mathcal{L}(\mathcal{D};\bm{\theta})}{\partial c_{k}}\right|_{\bm{\theta}_{\mathrm{MLE}}}$ $=-\sum_{h\in\Omega}2(\bar{C}^{h}-C^{h})v_{k}^{h}$ (S3)
$=-\sum_{h\in\Omega}2(b_{k}^{h}-v_{k}^{h}c_{k})v_{k}^{h},$
where $\mathcal{D}$ is the training dataset, $\bm{\theta}$ is the learnable model parameter vector, and we define $b_{k}^{h}:=\bar{C}^{h}-\sum_{j\neq k}v_{j}^{h}c_{j}$ to isolate the contribution of the $k$-th Gaussian to the rendered color at pixel $h$. Assuming a local minimum, the MLE estimate satisfies:
$\left.\frac{\partial\mathcal{L}(\mathcal{D};\bm{\theta})}{\partial c_{k}}\right|_{\theta_{\mathrm{MLE}}}=0$ (S4)
Solving this closed-form yields the optimal color for Gaussian $k$:
$c_{k}=\frac{\sum_{h\in\Omega}b_{k}^{h}v_{k}^{h}}{\sum_{h\in\Omega}(v_{k}^{h})^{2}}$ (S5)
Finally, by assuming $b_{k}^{h}\sim\mathcal{N}(v_{k}^{h}\bar{c}_{k},\epsilon^{2})$ where $\bar{c}_{k}$ is the groundtruth Gaussian color and $\mathrm{Var}(b_{k}^{h})=\epsilon^{2}$, the corresponding closed-form Gaussian color uncertainty (i.e., $\sigma_{k}^{2}:=\mathrm{Var}(c_{k}|v_{k}^{h})$) under the unit-variance Gaussian noise model (i.e., $\epsilon^{2}=1$) can be derived as:
$\sigma_{k}^{2}=\left(\sum_{h\in\Omega}(v_{k}^{h})^{2}\right)^{-1}$ (S6)
A.2 Details of Training Losses (Section 3 and Section 4.3 of the Main Paper)
As described in Section 4.3 and Eq. 14 of the main paper, our total loss is:
$\mathcal{L}^{\textrm{total}}=\mathcal{L}^{\textrm{rgb}}+\mathcal{L}^{\textrm{key}}+\mathcal{L}^{\textrm{non-key}}$ (S7)
where $\mathcal{L}^{\textrm{rgb}}$ is the perception loss and $\mathcal{L}^{\textrm{motion,key}}$ and $\mathcal{L}^{\textrm{motion,non-key}}$ are the motion loss used in $\mathcal{L}^{\textrm{key}}$ and $\mathcal{L}^{\textrm{non-key}}$, which covers the motion locality of key and non-key nodes, respectively. In this section, we provide additional details.
A.2.1 Motion Loss
The vanilla Gaussian splatting approach (e.g., Luiten et al. (2024)) represents a dynamic scene using 3D Gaussians with time-varying motion and define motion loss. The recent paper also follow up on the motion loss and modifies it based on their tasks. Here, we give the general form of these motion losses. The isometry loss describing the distance constraint between Gaussians is defined as
$\mathcal{L}^{\text{iso}}=\frac{1}{k|\mathcal{V}|}\sum_{t=0}^{T-1}\sum_{i\in\mathcal{V}}\sum_{j\in\text{knn}_{i;k}}w_{i,j}(\|\mathbf{p}_{j,n}-\mathbf{p}_{i,n}\|_{2}-\|\mathbf{p}_{j,t}-\mathbf{p}_{i,t}\|_{2})$ (S8)
where subscript $n$ means canonical space, $\mathbf{t}$ is the Gaussian’s 3D position vector, and $w_{i,j}$ represent the edge weight between $i$-th and $j$-th Gaussians. Since isometry does not take coordinates into consideration, we use rigidity loss to unify the coordinates and constrain the relative motion between Gaussians. The rigidity loss is defined by
$\mathcal{L}^{\text{rigid},\Delta}=\frac{1}{k|\mathcal{V}|}\sum_{t=\Delta}^{T-1}\sum_{i\in\mathcal{V}}\sum_{j\in\text{knn}_{i;k}}w_{i,j}\left\|\left(\mathbf{p}_{j,t-\Delta}-\mathbf{T}_{i,t-\Delta}\mathbf{T}_{i,t}^{-1}\mathbf{p}_{j,t}\right)\right\|_{2}$ (S9)
where $\mathbf{T}_{i,t}$ is the $i$-th Gaussian transformation matrix at $t$-th time index and $\Delta$ is the time interval. With a larger weight $w_{i,j}$, the Gaussian pair is more rigid. Beyond that, we also constrain the relative rotation explicitly for finer control on the rotation penalty. The relative rotation loss is defined by
$\mathcal{L}^{\text{rot},\Delta}=\frac{1}{k|\mathcal{V}|}\sum_{t=\Delta}^{T-1}\sum_{i\in\mathcal{V}}\sum_{j\in\text{knn}_{i;k}}w_{i,j}\left\|\mathbf{q}_{j,t}\mathbf{q}_{j,t-\Delta}^{-1}-\mathbf{q}_{i,t}\mathbf{q}_{i,t-\Delta}^{-1}\right\|_{2}$ (S10)
where $\mathbf{q}$ is the quaternion representation of the rotation matrix. All there three loss define the deformation of the Gaussians. Besides, the object motion in the world space tends to be smooth, so we define the velocity and acceleration to regularize the Gaussian model to avoid overfitting on the training views. The velocity loss is defined as
$\mathcal{L}^{\text{vel}}=\sum_{t=1}^{T-1}\sum_{i\in\mathcal{V}}\|\mathbf{p}_{i,t-\Delta}-\mathbf{p}_{i,t}\|_{1}+\|\mathbf{q}_{i,t-1}\mathbf{q}_{i,t}^{-1}\|_{1}$ (S11)
The acceleration loss is defined as
$\mathcal{L}^{\text{acc}}=\sum_{t=2}^{T-1}\sum_{i\in\mathcal{V}}\|\mathbf{p}_{i,t-2}-2\mathbf{p}_{i,t-1}+\mathbf{p}_{i,t}\|_{1}+\|\mathbf{q}_{i,t-2}\mathbf{q}_{i,t-1}^{-1}(\mathbf{q}_{i,t-1}\mathbf{q}_{i,t}^{-1})^{-1}\|_{1}$ (S12)
The motion loss used in the recent works (Lei et al., 2025; Stearns et al., 2024; Huang et al., 2024; Luiten et al., 2024) can be obtained by the different combinations of these losses. Then, the motion locality loss is defined as
$\mathcal{L}^{\text{motion}}=\lambda^{\text{iso}}\mathcal{L}^{\text{iso}}+\lambda^{\text{rigid}}\mathcal{L}^{\text{rigid}}+\lambda^{\text{rot}}\mathcal{L}^{\text{rot}}+\lambda^{\text{vel}}\mathcal{L}^{\text{vel}}+\lambda^{\text{acc}}\mathcal{L}^{\text{acc}},$ (S13)
where $\lambda$ is the hyperparameter. Specifically, we set $\lambda^{\mathrm{iso}}$ and $\lambda^{\mathrm{rigid}}$ to $1$ to ensure the geometry preservation and assign $\lambda^{\mathrm{rot}}$, $\lambda^{\mathrm{vel}}$, and $\lambda^{\mathrm{acc}}$ to $0.01$ for rigid orientation and motion smoothness.
A.2.2 Perception Loss
The perception loss $\mathcal{L}^{\text{rgb}}$ includes a standard combination of RGB $\ell_{1}$ loss and SSIM loss. Following the base models (SoM and MoSca), we also incorporate 2D prior losses, i.e., mask loss, depth loss, depth gradient loss, and tracking loss, into $\mathcal{L}^{\text{rgb}}$. We emphasize that these additional terms are not part of our proposed method, but are standard components inherited from the respective base models to ensure consistent training behavior.
Appendix B Implementation Details
USplat4D is compatible with dynamic Gaussian splatting methods that provide initial motion parameters. In this section, we describe implementation details for integrating USplat4D with two strong base models (i.e., SoM (Wang et al., 2025a) and MoSca (Lei et al., 2025)) along with our training and evaluation protocols.
B.1 Adaptation of SoM and MoSca
Both SoM (Wang et al., 2025a) and MoSca (Lei et al., 2025) adopt low-rank motion parameterizations to achieve compact, smooth, and rigid 4D representations. We first convert the outputs of SoM or MoSca into the data structure required by USplat4D. Specifically, we extract the 4D Gaussian primitives $\{\mathbf{G}_{i}\}_{i=1}^{N_{g}}$ from each method and reformat them into a unified representation that supports our graph construction and uncertainty modeling pipeline. To mitigate inconsistencies in Gaussian scale and distribution, we unify the spatial volume when selecting key Gaussians and normalize the uncertainty threshold used in both key node selection and edge construction. This ensures that our method behaves consistently across different base models. Since our method introduces additional optimization on top of the pretrained SoM and MoSca models, we ensure fair comparison by continuing to train the original SoM and MoSca with the same number of additional iterations as our baselines.
B.2 Training USplat4D Model
During preprocessing of USplat4D, we sequentially initialize the key and non-key graphs. To train the USplat4D model, we build on pretrained base models and allocate additional training iterations for fair comparison. Specifically, we train SoM for 400 extra epochs and MoSca for 1600 extra steps, using a consistent batch size of 8. These schedules are empirically chosen to ensure good convergence while aligning with the respective optimization routines of the base models.
For the first 10% and the last 20% of the training duration, we disable both density control and opacity reset to maintain stability. For the remaining iterations, we enable density control and opacity reset to improve rendering quality. The per-epoch and per-step training time is similar to that of SoM and MoSca, respectively.
Appendix C Additional Experimental Results
Table S1: Quantitative results on the DyCheck dataset (Gao et al., 2022). We compare our method applied to two dynamic Gaussian baselines (SoM and MoSca) with NeRF-based and Gaussian-based methods. We report results on 5 scenes at $1\times$ resolution and 7 scenes at $2\times$ resolution, following existing protocols. Our approach consistently improves the base models across all metrics and outperforms other baselines in both settings. Best results are in bold, second-best are underlined.
[…]
5 scenes
1 $\times$ resolution
DynIBaR (Li et al., 2023) 13.41 0.48 0.55
HyperNeRF (Park et al., 2021) 15.99 0.59 0.51
T-NeRF (Gao et al., 2022) 15.6 0.55 0.55
SC-GS (Huang et al., 2024) 14.13 0.477 0.49
Deformable 3DGS (Yang et al., 2024c) 11.92 0.490 0.66
4DGS (Wu et al., 2024) 13.42 0.490 0.56
MoDec-GS (Kwak et al., 2025) 15.01 0.493 0.44
MoBlender (Zhang et al., 2025) 16.79 0.650 0.37
HiMoR (Liang et al., 2025a) 0.46
SoM (Wang et al., 2025a) 16.72 0.630 0.45
USplat4D (ours) 16.85 0.650 0.38
7 scenes
2 $\times$ resolution
D-NPC (Kappel et al., ) 16.41 0.582 0.319
CTNeRF (Miao et al., 2024) 17.69 0.531
DyBluRF (Sun et al., 2024a) 17.37 0.591 0.373
HyperNeRF (Park et al., 2021) 16.81 0.569 0.332
T-NeRF (Gao et al., 2022) 16.96 0.577 0.379
PGDVS (Zhao et al., 2024) 15.88 0.548 0.34
Dynamic Gaussians (Luiten et al., 2024)   7.29 0.69
4DGS (Wu et al., 2024) 13.64 0.43
Gaussian Marbles (Stearns et al., 2024) 16.72 0.41
MoSca (Lei et al., 2025) 19.32 0.706 0.26
USplat4D (ours) 19.63 0.716 0.25
C.1 Comparison with Additional NeRF-based Methods on DyCheck Dataset
In Table 1 of the main paper, we primarily compare with Gaussian Splatting based methods that use explicit 3D representations. Here, we extend the comparison by including additional NeRF-based methods (Miao et al., 2024; Park et al., 2021; Gao et al., 2022; Li et al., 2023; Sun et al., 2024a; Kappel et al., ; Zhao et al., 2024) that adopt implicit neural radiance fields. While these methods are effective for static or mildly dynamic scenes, our experiments show that Gaussian splatting based approaches consistently outperform NeRF-based methods on the DyCheck dataset (Gao et al., 2022). This comparison serves as an extension of Table 1 and the results are summarized in Table S1.
C.2 Tracking Results on the DyCheck Dataset
[…]
Table S2: Tracking results on the DyCheck dataset (Gao et al., 2022). We follow the evaluation protocols of MoSca and SoM to report 3D keypoint tracking metrics. USplat4D outperforms both baselines by a large margin. Please refer to our supplementary video demo for qualitative results.
[…]
MoSca (Lei et al., 2025) 82.4 SoM (Wang et al., 2025a) 0.082 43.0 73.3
USplat4D (ours) 84.5 USplat4D (ours) 0.072 54.4 75.8
C.3 Per-Scene Results on the DyCheck Dataset
We further present a per-scene breakdown of the DyCheck results reported in Table 1 of the main paper. Results are shown in Table S3. Our method consistently improves per-scene performance across most scenes across PSNR, SSIM, and LPIPS.
Table S3: Per-scene results on the DyCheck dataset (Gao et al., 2022). We provide a detailed breakdown of the results summarized in Table 1 of the main paper. $\star$: results reproduced by us, as the original numbers in Table 1 were directly reported from the SoM and MoSca papers. Minor discrepancies may exist due to differences in training. We note that the validation views are near the training views in the DyCheck dataset. Therefore, the observed improvements do not fully reflect the advantages of USplat4D in extreme novel view synthesis. Results are formatted as PSNR ($\uparrow$) / SSIM ($\uparrow$) / LPIPS ($\downarrow$).
(a) 7 Scenes, 2 $\times$ Resolution
Method Apple Block Spin Paper Windmill
4DGS (Wu et al., 2024) 15.41 / 0.450 / – 11.28 / 0.633 / – 14.42 / 0.339 / – 15.60 / 0.297 / –
Gaussian Marbles 17.70 / 0.492 / – 17.42 / 0.384 / – 18.88 / 0.428 / – 17.04 / 0.394 / –
[…]
USplat4D (ours) 19.53 / 0.813 / 0.32 18.49 / 0.681 / 0.31 21.77 / 0.772 / 0.16 22.55 / 0.753 / 0.14
Method Space-Out Teddy Wheel Mean
Gaussian Marbles 15.94 / 0.435 / – 13.95 / 0.442 / – 16.14 / 0.351 / – 16.72 / 0.418 / –
[…]
USplat4D (ours) 20.81 / 0.668 / 0.24 15.78 / 0.625 / 0.36 18.46 / 0.696 / 0.22 19.63 / 0.716 / 0.25
(b) 5 Scenes, 1 $\times$ Resolution
Method Apple Block Spin
SC-GS (Huang et al., 2024) 14.96 / 0.692 / 0.51 13.98 / 0.548 / 0.48 14.32 / 0.407 / 0.45
MoDec-GS (Kwak et al., 2025) 16.48 / 0.699 / 0.40 15.57 / 0.590 / 0.48 15.53 / 0.433 / 0.37
[…]
USplat4D (ours) 16.94 / 0.754 / 0.49 16.12 / 0.653 / 0.45 17.75 / 0.711 / 0.27
Method Paper Windmill Teddy Mean
SC-GS (Huang et al., 2024) 14.87 / 0.221 / 0.43 12.51 / 0.516 / 0.56 14.13 / 0.477 / 0.49
MoDec-GS (Kwak et al., 2025) 14.92 / 0.220 / 0.38 12.56 / 0.521 / 0.60 15.01 / 0.493 / 0.44
[…]
USplat4D (ours) 19.69 / 0.555 / 0.19 13.77 / 0.551 / 0.50 16.85 / 0.645 / 0.38
C.4 Data Curation for the Objaverse Dataset
To evaluate our model under large view shifts, which are rarely present in existing realistic datasets such as DyCheck (Gao et al., 2022) and DAVIS (Perazzi et al., 2016), we curate a high-quality dynamic 3D dataset sourced from Objaverse-1.0 (Deitke et al., 2023). Each curated scene contains a single object with one or more animation labels; by default, we animate the object using its first animation label. As described in Section 5.1 of the main paper, we render 121 frames for each view. Because the total number of animation frames in Objaverse is typically smaller than 121, we cycle through the animation frames in a zigzag pattern to maintain motion continuity across the rendered sequence. We place 12 cameras evenly around the object at 30° horizontal intervals and move them along a circular trajectory in a counterclockwise direction with a 3° step per frame. For each frame, we render RGB, depth, and mask images using Blender (Blender Online Community, 2018) and record the corresponding camera intrinsics and extrinsics as ground truth. In our experiments, we use one of the 12 monocular videos as input and utilize the left 11 monocular videos to evaluate rendering quality over three angular ranges: $(0^{\circ},60^{\circ}]$, $(60^{\circ},120^{\circ}]$, and $(120^{\circ},180^{\circ}]$.
C.5 Additional Results on the Objaverse Dataset
Figure S1: Additional qualitative results on Objaverse. Each case shows a 3D rendering from an input view and a comparison between the baseline (SoM (Wang et al., 2025a) or MoSca (Lei et al., 2025)) and our USplat4D at an extreme novel view (red). Please see the supplementary video for clearer visualization.
Table S4: Per-scene results on the Objaverse dataset. We provide a detailed breakdown of the results summarized in Table 2 of the main paper. We evaluate novel view synthesis across increasing horizontal angular ranges: $(0^{\circ},60^{\circ}]$, $(60^{\circ},120^{\circ}]$, and $(120^{\circ},180^{\circ}]$.
Scene Method View Range $(0^{\circ},60^{\circ}]$ View Range $(60^{\circ},120^{\circ}]$ View Range $(120^{\circ},180^{\circ}]$
[…]
Chick SoM 14.26 0.852 0.34 13.86 0.845 0.37 14.70 0.847 0.36
USplat4D 14.59 0.852 0.32 14.60 0.853 0.32 15.13 0.857 0.31
MoSca 14.69 0.869 0.27 14.35 0.864 0.29 14.47 0.867 0.29
USplat4D 14.82 0.880 0.24 14.42 0.876 0.27 14.97 0.878 0.25
Panda SoM 15.15 0.857 0.30 14.51 0.843 0.33 15.06 0.852 0.31
USplat4D 15.51 0.873 0.26 15.28 0.874 0.26 15.38 0.874 0.26
MoSca 15.37 0.888 0.21 14.83 0.881 0.23 15.15 0.880 0.22
USplat4D 15.08 0.888 0.22 14.33 0.883 0.24 15.03 0.886 0.22
Woman SoM 18.47 0.900 0.21 17.92 0.894 0.22 19.85 0.904 0.20
USplat4D 20.34 0.920 0.13 20.28 0.920 0.13 21.07 0.923 0.12
MoSca 16.67 0.910 0.15 16.48 0.908 0.15 16.79 0.910 0.14
USplat4D 17.00 0.910 0.13 17.07 0.911 0.13 17.16 0.911 0.12
Rat SoM 16.61 0.833 0.36 15.81 0.837 0.38 16.08 0.835 0.37
USplat4D 16.73 0.833 0.35 16.66 0.837 0.35 17.04 0.835 0.34
MoSca 17.94 0.859 0.27 17.19 0.848 0.29 16.85 0.847 0.30
USplat4D 18.53 0.873 0.23 18.09 0.872 0.24 18.27 0.874 0.24
Doctor SoM 17.78 0.877 0.28 17.51 0.874 0.29 18.18 0.876 0.28
USplat4D 17.68 0.885 0.25 17.35 0.885 0.26 17.77 0.887 0.25
MoSca 17.12 0.896 0.23 16.16 0.887 0.26 16.76 0.888 0.24
USplat4D 17.07 0.898 0.21 16.32 0.894 0.24 17.10 0.897 0.22
Angry Bird SoM 14.23 0.839 0.34 13.85 0.831 0.34 14.84 0.836 0.32
USplat4D 15.24 0.843 0.32 15.27 0.845 0.30 15.79 0.855 0.29
MoSca 15.04 0.864 0.29 15.16 0.859 0.28 14.97 0.859 0.29
USplat4D 14.78 0.864 0.28 14.60 0.862 0.28 15.07 0.868 0.26
Along with the qualitative results in Figure 6 of the main draft, three additional cases are provided in Figure S1 for further comparison. We further provide per-scene results for the Objaverse dataset in Table S4, complementing the summary reported in Table 2 of the main draft. Our method consistently achieves lower LPIPS than both SoM (Wang et al., 2025a) and MoSca (Lei et al., 2025) across all angular ranges and scenes. For PSNR and SSIM, our model is also generally superior to SoM and MoSca in most angular ranges and scenes.
C.6 Results on NVIDIA Dataset
Our main quantitative analysis focuses on the DyCheck-iPhone dataset (Gao et al., 2022) in the main paper, which aligns well with our goal of synthesizing dynamic objects captured by a moving camera (see Section 1 of the main paper). Here we also evaluate our method on the NVIDIA dataset (Yoon et al., 2020) for broader comparison by following the evaluation protocal in RDRF (Liu et al., 2023). The results are shown in Table S5.
We note that USplat4D also improves performance when applied to state-of-the-art MoSca (Lei et al., 2025) on the NVIDIA dataset, although the gains are marginal. This is expected, as the NVIDIA dataset (or other datasets such as HyperNeRF (Park et al., 2021) dataset) differs from DyCheck to feature input views with more limited motion.
Table S5: Results on the NVIDIA dataset Yoon et al. (2020).
[…]
MoSca Lei et al. (2025) 26.77 0.854 0.07
MoSca + G2DSplat 26.93 0.855 0.07
Appendix D Additional Ablation and Analysis
D.1 Ablation Study on Key / Non-Key Ratio
We select the uncertainty threshold by keeping the top 2% (1000-th) of all Gaussians. This strategy ensures broad spatial coverage while filtering unreliable nodes. We find this setting to be robust across scenes and base models. To further evaluate the impact of the threshold, we conduct an ablation study with different thresholds. As shown in Table S6, the results remain stable, and our method is not sensitive to this parameter. We vary the ratio from 0.005 to 0.04 and observed consistent performance, including our chosen value of 0.02, which demonstrates that using this well-justified parameter is reasonable and effective.
Table S6: Ablation study on the uncertainty thresholds. Results are reported relative to our chosen ratio threshold (0.02).
[…]
MoSca (baseline) 0.994 0.995 1.024
0.005 1.000 1.000 1.000
0.01 1.001 1.001 0.999
0.02 (ours, reference) 1.000 1.000 1.000
0.03 1.001 1.001 0.998
0.04 1.001 1.000 1.000
D.2 Analysis on Key Node Selection
In Section 4.2, we design the key node selection strategy by sampling key node candidates via 3D gridization to ensures spatial coverage and reduces redundancy. To test its functionality, we replace it by a random sampling strategy. Specifically, we keep the number of selected Gaussians per frame the same, but sample them randomly from all remaining grids without enforcing per-grid constraints. Table S7 demonstrates the performance drops on the DyCheck validation set, which causes by a non-uniform spatial distribution (i.e., some grids may contain multiple selected nodes, while some are missed entirely).
Table S7: Ablation study on key node selection strategy. We assess the impact of the key node sampling strategy by randomly sampling without considering spatial distribution.
[…]
USplat4D (full model) 19.63 0.716 0.249
   w/o spatial-random key node selection 19.50 0.712 0.251
D.3 Analysis on Runtime
[…]
Table S8: Ablation for comparable time on DyCheck dataset.
[…]
MoSca (baseline) 19.32 0.706 0.264
Ours (same time with MoSca) 19.41 0.710 0.254
Ours (full time) 19.63 0.715 0.249
D.4 Limitations and Failure Cases
Although the model partially relies on the tracking results of the initial 4D Gaussians and the geometry information during the observation time period, both quantitative and qualitative results demonstrate the robustness of the model in restoring or preserving shape when faced with partial observation and tracking failure. However, the graph model still suffers from the incorrect motion in the 4D prior. Although our model can utilize uncertainty model to distinguish the usefully information but if the 4D prior totally fails without leaving useful information, our model can not retrieve useful information from it. According to our observations, when the video only includes few frames with large motion and the initial 4D Gaussians do not provide stable tracking results, the output of the USplat4D may not guarantee the geometry consistency. The failure case can be found in the demo video.
Appendix E Social and Broader Impact
Our work advances dynamic 3D scene reconstruction and novel view synthesis from monocular videos, with applications in AR/VR, physical scene understanding, digital content creation, and human-object interaction. By modeling uncertainty and improving synthesis under extreme viewpoints, our method contributes to more robust and accessible 4D modeling. While the method has positive use cases, it could be misused for synthetic content manipulation. Our method does not involve sensitive data, and care should be taken in downstream applications to ensure responsible use.